# Magic Clothing: Controllable Garment-Driven Image Synthesis

### v1

@@ Abstract.
ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer vision problemsccs: Computing methodologies Supervised learning
@@ 2.1. Latent Diffusion Models
Latent diffusion models (LDMs) (Rombach et al., 2022) have been successfully used for text-to-image generation tasks. Based on this robust foundational model, various LDM-based researches and applications have emerged. To add spatial conditioning controls to the pretrained LDMs, ControlNet (Zhang et al., 2023b) incorporated trainable encoder blocks into the original UNet. Meanwhile, T2I-Adapter (Mou et al., 2023) proposed a compact network design that provided the same functionality as ControlNet but with reduced complexity. To further reduce the training cost, Uni-ControlNet (Zhao et al., 2024b) proposed a unified framework that handles different conditional controls in a flexible and composable manner within one single model. On the other hand, LDMs have also played a significant role in the domain of image editing. InstructPix2Pix (Brooks et al., 2023) retrained the UNet of LDMs by adding extra input channels to the first convolutional layer on a large dataset of image editing examples to make it follow the edit instructions. MasaCtrl (Cao et al., 2023) converted the self-attention in diffusion models into mutual self-attention to enable consistent image generation and complex non-rigid image editing simultaneously without additional training cost. InfEdit (Xu et al., 2023a) performed consistent and faithful image editing for both rigid and non-rigid semantic changes through the denoising diffusion consistent model and attention control mechanisms. In this paper, we make full use of the power of pretrained LDMs in text-to-image generation, and introduce an extra garment extractor for garment-driven image synthesis.
@@ 2.2. Subject-Driven Image Synthesis
Given an image of a particular subject, the subject-driven image synthesis task is to generate a novel image in different contexts while maintaining high fidelity to its key visual features. Textual Inversion (Gal et al., 2022) learned to represent the concept through new ”words” in the embedding space of a frozen text-to-image model, which can be composed into natural language sentences and generate customized images. DreamBooth (Ruiz et al., 2023a) finetuned the pretrained LDMs model with the newly designed class-specific prior preservation loss to make it learn to bind a unique identifier with that specific subject. However, finetuning the entire UNet for each subject raises significant computational cost. To address this, HyperDreamBooth (Ruiz et al., 2023b) proposed a hypernetwork capable of efficiently generating an initial prediction of a subset of network weights and significantly accelerated the finetuning process. BLIP-Diffusion (Li et al., 2024) introduced a new multimodal encoder which is pretrained to provide subject representation. By leveraging such visual representation, diffusion models can yield a wide range of subject variations. IP-Adapter (Ye et al., 2023) designed a decoupled cross-attention mechanism that separates cross-attention layers for text features and image features. Despite its simplicity, this method achieved comparable or even better performance than specifically finetuned models. Versatile Diffusion (Xu et al., 2023b) expanded the existing single-flow diffusion pipeline into a multitask multimodal network to enable cross-modal generality. Break-A-Scene (Avrahami et al., 2023) extracted a distinct text token for each concept from a single image by introducing a two-phase customization process that optimizes a set of dedicated textual embedding and the model weights. In this work, we focus on an unexplored garment-driven image synthesis task, which requires much better detail preservation together with the creative capability following diverse text prompts.
@@ 3.2. Network Architecture
(2) ${\rm{Attention}}(\alpha_{i},\beta_{i})={\rm{Softmax}}(\frac{\mathbf{W}_{Q}\beta_{i}(\mathbf{W}_{K}[\alpha_{i},\beta_{i}])^{T}}{\sqrt{d}})\mathbf{W}_{V}[\alpha_{i},\beta_{i}],$
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation operation, $d$ is the feature dimension, and $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ are linear projection weights of query, key, and value in self-attention layers.
@@ 3.3. Joint Classifier-free Guidance
Classifier-free guidance (Ho and Salimans, 2022) is a method that helps the diffusion models to attain a trade-off between sample quality and diversity by jointly training a conditional and an unconditional diffusion model. The implementation during training is relatively simple by setting the conditional control $\mathbf{c}=\varnothing$ with some probability. During inference, the score estimate $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c})$ linearly combine the conditional and unconditional predictions as the following equation:
@@ 4.1.3. Evaluation Metrics
To evaluate the garment fidelity, the details of the garment on the generated character must be compared with those of the original garment. Following previous subject-driven methods (Ruiz et al., 2023a, b), we use CLIP-I (Radford et al., 2021) to compute the similarity of the same subject in different images. However, this metric is sensitive to background and the character pose in our task. Recently, S-LPIPS (Zhang et al., 2024) is designed to tackle the problem of changes in the shape and posture of generated characters, but it still relies on human pose estimation (Wei et al., 2016; Simon et al., 2017; Cao et al., 2017) and paired images. Surprisingly, it is reported in (Tang et al., 2023) that geometric corresponding points can be accurately predicted from the intermediate layer activations of pretrained LDMs, even though they are not trained with any explicit geometry supervision. Driven by the above prospects and challenges, we propose Matched-Points-LPIPS (MP-LPIPS) that measures the LPIPS distance (Zhang et al., 2018) between pixels on garment $\mathbf{I}_{G}$ and matched pixels on generated character $\mathbf{I}_{C}$. Specifically, we uniformly select N points $\mathbb{P}_{G}$ on each garment, calculating their diffusion features (Tang et al., 2023) $\mathcal{F}_{\mathbf{t,l}}(\mathbb{P}_{G})$ at time step $t$ and layer $l$. Then we find the matched points $\mathbb{P}_{C}$ on the target character by calculating cosine similarity as:
@@ 5. Conclusion
This paper presents Magic Clothing, an LDM-based network architecture for the unexplored garment-driven image synthesis task. A garment extractor is employed to incorporate the garment details via self-attention fusion, and the joint classifier-free guidance is applied to balance the control of garment features and text prompts. Our garment extractor is a plug-in module which can be easily combined with other useful extensions for additional conditions. Comprehensive experiments demonstrate our superiority in generating diverse images controlled by the given garments and text prompts, implying the tremendous potential of controllable garment-driven image synthesis.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiadoi: 10.1145/3664647.3680691isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer vision problemsccs: Computing methodologies Supervised learning
@@ 2.1. Latent Diffusion Models
Latent diffusion models (LDMs) (Rombach et al., 2022) have been successfully used for text-to-image generation tasks. To add spatial conditioning controls to the pretrained LDMs, ControlNet (Zhang et al., 2023b) incorporated trainable encoder blocks into the original UNet. Meanwhile, T2I-Adapter (Mou et al., 2023) proposed a compact network design that provided the same functionality as ControlNet but with reduced complexity. To further reduce the training cost, Uni-ControlNet (Zhao et al., 2024b) proposed a unified framework that handles different conditional controls in a flexible and composable manner within one single model. On the other hand, LDMs have also played a significant role in the domain of image editing. InstructPix2Pix (Brooks et al., 2023) retrained the UNet of LDMs by adding extra input channels to the first convolutional layer on a large dataset of image editing examples to make it follow the edit instructions. MasaCtrl (Cao et al., 2023) converted the self-attention in diffusion models into mutual self-attention to enable consistent image generation and complex non-rigid image editing simultaneously without additional training cost. InfEdit (Xu et al., 2023a) performed consistent and faithful image editing for both rigid and non-rigid semantic changes through the denoising diffusion consistent model and attention control mechanisms. In this paper, we make full use of the power of pretrained LDMs in text-to-image generation, and introduce an extra garment extractor for garment-driven image synthesis.
@@ 2.2. Subject-Driven Image Synthesis
Given an image of a particular subject, the subject-driven image synthesis task is to generate a novel image in different contexts while maintaining high fidelity to its key visual features. Textual Inversion (Gal et al., 2022) learned to represent the concept through new ”words” in the embedding space of a frozen text-to-image model, which can be composed into natural language sentences and generate customized images. DreamBooth (Ruiz et al., 2023a) finetuned the pretrained LDMs model with the newly designed class-specific prior preservation loss to make it learn to bind a unique identifier with that specific subject. However, finetuning the entire UNet for each subject raises significant computational cost. To address this, HyperDreamBooth (Ruiz et al., 2023b) proposed a hypernetwork capable of efficiently generating an initial prediction of a subset of network weights and significantly accelerated the finetuning process. BLIP-Diffusion (Li et al., 2024) introduced a new multimodal encoder pretrained to provide subject representation. By leveraging such visual representation, diffusion models can yield a wide range of subject variations. IP-Adapter (Ye et al., 2023) designed a decoupled cross-attention mechanism that separates cross-attention layers for text features and image features. Despite its simplicity, this method achieved comparable or even better performance than specifically finetuned models. Versatile Diffusion (Xu et al., 2023b) expanded the existing single-flow diffusion pipeline into a multitask multimodal network to enable cross-modal generality. Break-A-Scene (Avrahami et al., 2023) extracted a distinct text token for each concept from a single image by introducing a two-phase customization process that optimizes a set of dedicated textual embedding and the model weights. In this work, we focus on an unexplored garment-driven image synthesis task, which requires much better detail preservation together with the creative capability following diverse text prompts.
@@ 3.2. Network Architecture
(2) $\begin{split}{\rm{Attention}}(\alpha_{i},\beta_{i})=&{\rm{Softmax}}(\frac{\mathbf{W}_{Q}\alpha_{i}(\mathbf{W}_{K}\alpha_{i})^{T}}{\sqrt{d}})\mathbf{W}_{V}\alpha_{i}+\\ &{\rm{Softmax}}(\frac{\mathbf{W}_{Q}\alpha_{i}(\mathbf{W}_{K_{G}}\beta_{i})^{T}}{\sqrt{d}})\mathbf{W}_{V_{G}}\beta_{i},\end{split}$
where $d$ is the feature dimension, $\mathbf{W}_{Q},\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ are frozen linear projection weights of query, key and value in self-attention layers. $\mathbf{W}_{K_{G}}$ and $\mathbf{W}_{V_{G}}$ are additional trainable projection weights of key and value for garment feature $\beta_{i}$.
@@ 3.3. Joint Classifier-free Guidance
Classifier-free guidance (Ho and Salimans, 2022) is a method that helps the diffusion models to attain a trade-off between sample quality and diversity by jointly training a conditional and an unconditional diffusion model. The implementation during training is relatively simple by setting the conditional control $\mathbf{c}=\varnothing$ with some probability. During inference, the score estimate $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{c})$ linearly combine the conditional and unconditional predictions as:
@@ 4.1.3. Evaluation Metrics
To evaluate the garment fidelity, the details of the garment on the generated character must be compared with those of the original garment. Following previous subject-driven methods (Ruiz et al., 2023a, b), we use CLIP-I (Radford et al., 2021) to compute the similarity of the same subject in different images. However, this metric is sensitive to background and the character pose in our task. Recently, S-LPIPS (Zhang et al., 2024) is designed to tackle the problem of changes in the shape and posture of generated characters, but it still relies on human pose estimation (Wei et al., 2016; Simon et al., 2017; Cao et al., 2017) and paired images. Surprisingly, it is reported in (Tang et al., 2023) that geometric corresponding points can be accurately predicted from the intermediate layer activations of pretrained LDMs, even though they are not trained with any explicit geometry supervision. Driven by the above prospects and challenges, we propose Matched-Points-LPIPS (MP-LPIPS) that measures the LPIPS distance (Zhang et al., 2018) between pixels on garment $\mathbf{I}_{G}$ and matched pixels on generated character $\mathbf{I}_{C}$. Specifically, we uniformly select N points $\mathbb{P}_{G}$ on each garment, calculating their diffusion features (Tang et al., 2023) $\mathcal{F}_{\mathbf{t,l}}(\mathbb{P}_{G})$ at time step $t$ and layer $l$. Then we find the matched points $\mathbb{P}_{C}$ on the target character by calculating the cosine similarity as:
@@ 5. Conclusion
This paper presents Magic Clothing, an LDM-based network architecture for the unexplored garment-driven image synthesis task. A garment extractor is employed to incorporate the garment details via self-attention fusion, and the joint classifier-free guidance is applied to balance the control of garment features and text prompts. Our garment extractor is a plug-in module which can be easily combined with other useful extensions for additional conditions. Comprehensive experiments demonstrate our superiority in generating diverse images controlled by the given garments and text prompts, implying the potential of garment-driven image synthesis.
Acknowledgements.
To Eve Yan and Carol Shu, for their beautiful portraits.
Appendix A MP-LPIPS details
Table 3 provides the detailed parameter settings for Matched-Points-LPIPS (MP-LPIPS), which are omitted in the paper for brevity. We first uniformly sample points $\mathbb{P}_{G}$ inside the garment area of the garment image $\mathbf{I}_{G}$ given the garment mask from the VITON-HD test dataset (Choi et al., 2021) to ensure the sample distance $d_{s}=40$ pixels between each point and its nearest neighbour. For sampled points $\mathbb{P}_{G}$ on $\mathbf{I}_{G}$, we retrieve their corresponding matching points $\mathbb{P}_{C}$ on the target character $\mathbf{I}_{C}$ using the diffusion features (Tang et al., 2023) at time step $t=41$ and $11$-th layer inside UNet upsampling blocks. We eventually calculate MP-LPIPS as the average LPIPS distance of the patches $\mathbf{P}_{G},\mathbf{P}_{C}$ centred on matched points $\mathbb{P}_{G}$ and $\mathbb{P}_{C}$, where the patch size $s=33\times 33$ pixels and we black out the area except for the garment part. If any of $\mathbf{P}_{C}$ fall outside the garment area of $\mathbf{I}_{C}$ by more than $\tau_{s}=17$ pixels, we view those points as mismatched and set their LPIPS to $0.6$ as the mismatch penalty. Here the garment area of $\mathbf{I}_{C}$ is obtained by applying HumanParsing (Li et al., 2020) on $\mathbf{I}_{C}$.
Name Value
time step $t$ 41
layer $l$ 11
patch size $s$ $33\times 3$ 33
sample distance $d_{s}$ 40
sample threshold $\tau_{s}$ 17
mismatch penalty $p$ 0.6
Table 3. Detailed Settings for the MP-LPIPS.
Appendix B Failure Cases
Figure 7. Failure cases of our Magic Clothing. Based on Realistic Vision V4.0, it fails to generate an anime-style character (1st row left), and vice versa for Counterfeit V3.0 (1st row right). Example results of complicated garments like down jackets and coats are shown in the second row.
In Figure 7 we present several failure cases as discussed in the paper. While our model achieves remarkable results in garment-driven image synthesis and is compatible with various finetuned LDMs, the styles of generated images are highly dependent on the base diffusion models. In practice, we apply our model to Stable Diffusion V1.5 (Rombach et al., 2022) finetuned on the photorealistic data (i.e., Realistic Vision V4.0111https://huggingface.co/SG161222) and Stable Diffusion V1.5 finetuned on the anime-style data (i.e., Counterfeit V3.0222https://huggingface.co/gsdf/Counterfeit-V3.0), respectively. As shown in the first row of Figure 7, our model tends to generate realistic characters with Realistic Vision V4.0 and anime-style characters with Counterfeit V3.0, regardless of the target styles specified by the text prompts. Another limitation is that due to limited garment types of training samples in the VITON-HD dataset, our model may not perform perfectly on complicated garments. As can be seen in the second row of figure 7, our model fails to preserve the middle layer of the given down jacket and slightly reduces the size of the overcoat.